# InvertAvatar: Incremental GAN Inversion for Generalized Head Avatars

### v2

@@ 6.2.3 Training $E_{tri\_rec}$ and $E_{tex\_rec}$
We adopt the same loss terms and training strategy as in Sec 6.2.2. On eight 3090 GPUs with a batch size of 8, we train both encoders without the adversarial objective for about 3 days, and then finetune with adversarial loss for additional 24 hours.
6.3 Network Architecture
@@
To implement the latent encoder $E_{latent}$, we adopt the design of the e4e encoder [45] and employ IR-SE-50 [10] pretrained model for the backbone network.
@@
Adopting conventional U-Net architecture, $E_{tri}$ and $E_{tex}$ are designed based on TriplaneNet [6]. As both networks produce multi-resolution feature maps, we incorporate an extra MLP block at each layer to transform the output channels.
@@
7 More Qualitative Results
In Fig. 9, Fig. 10, Fig. 11, Fig. 13 and Fig. 12, we present more visualizations of cross-identity reenactment. Please also refer to our video for more clear comparisons.
Figure 9: Cross-identity reenactment results of one-shot avatar reconstruction (part 1/3).
Figure 10: Cross-identity reenactment results of one-shot avatar reconstruction (part 2/3).
Figure 11: Cross-identity reenactment results of one-shot avatar reconstruction (part 3/3).
Figure 12: Cross-identity reenactment results of few-shot avatar reconstruction.
Figure 13: Cross-identity reenactment results of one-shot avatar reconstruction. We illustrate the multi-view rendering results and geometry.

### v3

@@ A.2.3. Training $E_{tri\_rec}$ and $E_{tex\_rec}$
We adopt the same loss terms and training strategy as in Sec A.2.2. On eight 3090 GPUs with a batch size of 8, we train both encoders without the adversarial objective for about 3 days, and then finetune with adversarial loss for additional 24 hours.
A.3. Network Architecture
@@
To implement the latent encoder $E_{latent}$, we adopt the design of the e4e encoder (Tov et al., 2021) and employ IR-SE-50 (Deng et al., 2019) pretrained model for the backbone network.
@@
Adopting conventional U-Net architecture, $E_{tri}$ and $E_{tex}$ are designed based on TriplaneNet (Bhattarai et al., 2023). As both networks produce multi-resolution feature maps, we incorporate an extra MLP block at each layer to transform the output channels.
@@
Figure 12. LPIPS over the number of input frames on VFHQ-test. We demonstrate the effectiveness of our method to handle unordered input. Besides, we conduct a comparison with an Attention-model baseline to validate our choice of recurrent network.
Appendix B More Experiments
B.1. Reason for the choice of Recurrent Network
We choose RNN for two reasons. Firstly, for few-shot tasks, the length of the input image sequence usually does not exceed 50, and GRU can handle such short sequences without suffering from forgetting issues. Secondly, despite Attention’s widespread use for unordered sequences, its high computational cost limits supported sequence lengths during training. In contrast, GRU’s computation grows linearly with sequence length, allowing the utilization of more frames for learning long-term temporal aggregation during training.
To evaluate our recurrent mechanism, we establish a fixed-window baseline termed “AttFusion”, which replaces each GRU block with a transformer block to merge temporal features. As we can only train it with 3 frames in our practice, during inference “AttFusion” averages the inverted canonical features when handling longer sequences.
To demonstrate our ability to handle non-sequential input, we randomly sample from a video to form input sequences. It can be observed that the two curves representing our method in Fig. 12 and Fig. 6 almost overlap, indicating that our approach is comparably effective in handling both ordered and unordered data.
As illustrated in Fig. 12, Our ConvGRU-based implementation and Attention models have similar performance for short sequences. However, for longer sequences, the performance of the Attention baseline suffers from the blur caused by the averaging strategy, and our recurrent framework’s performance improves incrementally with additional frames. Hence, we think ConvGRU is more suitable for our flexible few-shot task.
B.2. Ablation on the network design of Animatable 3D GAN
As mentioned in Sec. 3.2 and Sec. A.1, we replace the two UNets in Next3D with a single generator $G_{face}$. To evaluate the impact of the replacement, we use the FFHQ dataset to train Next3D network with the same facial parametric model (FaceVerse), and illustrate the quantitative comparison in Tab. 5.
Despite a slight increase in FID score due to fewer parameters, the improvement in AKD score demonstrates that our generator has more precise expression-driving performance. This is attributed to $G_{face}$’s multi-scale conditioning, which enables the conditional parametric model to directly control generated facial features. In contrast, the two cascaded post-processing UNets diminish control sensitivity.
[…]
Ours 4.1 0.057 0.853 84.21M
Next3D (FV) 3.9 0.095 0.852 164.81M
Table 5. Quantitative comparison with Next3D (FaceVerse) on image synthesis quality, animation accuracy, identity preservation and memory consumption.
Figure 13. Cross-identity reenactment results of one-shot avatar reconstruction. We illustrate the multi-view rendering results and geometry.